one important factor in the total delay of a store-and-forward packet-switching network is how long it takes to store and forward a packet through a switch. If the switching time of each switching device is 1 microsecond, is this likely to be a major factor in the response of a client-server system where the client is a computer you use to access the internet, and the server is somewhere in Europe? Explain your answer. You can assume the average propagation speed in all media involved is 2/3 the speed of light in a vacuum; you may calculate or estimate the distance between the server and your computer; you need to find out the number of routers/switches in between using Traceroute or Tracert.
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Speed of light in vacuum is 300,000 km/sec

Speed in copper is (2/3)*300,000=200,000

Switching time is 1 microsec

So,

At 0.000001 sec; it will travel to 0.2 km

Assume there are 50 switches. So, it adds to 10 km

If distance between New York and California is 5,000 km, the percentage of the equivalent
added distance is =1000/5000=0.2%

Kindly revert for any queries

Thanks.